Solve for x. Round to the nearest ten-thousandth.
5^(-1x+1) = 15

Answer by Alan3354(69443)

5^(-1x+1) = 15
(1-x)*log(5) = log(15)
log(5) - x*log(5) = log(15)
x*log(5) = log(5) - log(15) = log(1/3)
x = log(1/3)/log(5)
x =~ -0.6826